\begin{thebibliography}{10}

\bibitem{Aubin-MR0448404}
Thierry Aubin.
\newblock Probl\`emes isop\'erim\'etriques et espaces de {S}obolev.
\newblock {\em J. Differential Geometry}, 11(4):573--598, 1976.

\bibitem{Aubin2-MR534672}
Thierry Aubin.
\newblock Meilleures constantes dans le th\'eor\`eme d'inclusion de {S}obolev
  et un th\'eor\`eme de {F}redholm non lin\'eaire pour la transformation
  conforme de la courbure scalaire.
\newblock {\em J. Funct. Anal.}, 32(2):148--174, 1979.

\bibitem{BartolucciChenLinTarantello-MR2071339}
Daniele Bartolucci, Chiun-Chuan Chen, Chang-Shou Lin, and Gabriella Tarantello.
\newblock Profile of blow up solutions to mean field equations with singular
  data.
\newblock {\em Comm. Partial Differential Equations}, 29(7--8):1241--1265,
  2004.

\bibitem{BD-MR3355004}
Daniele Bartolucci and Francesca De~Marchis.
\newblock Supercritical mean field equations on convex domains and the
  {O}nsager's statistical description of two-dimensional turbulence.
\newblock {\em Arch. Ration. Mech. Anal.}, 217(2):525--570, 2015.

\bibitem{BartolucciGuiJevnikarMoradifam-MR3961327}
Daniele Bartolucci, Changfeng Gui, Alessio Jevnikar, and Amir Moradifam.
\newblock A singular sphere covering inequality: uniqueness and symmetry of
  solutions to singular liouville-type equations.
\newblock {\em Math. Ann.}, 374(3--4):1883--1922, 2019.

\bibitem{BartolucciLin-MR2542689}
Daniele Bartolucci and Chang-Shou Lin.
\newblock Uniqueness results for mean field equations with singular data.
\newblock {\em Comm. Partial Differential Equations}, 34(7):676--702, 2009.

\bibitem{BLT-MR2838340}
Daniele Bartolucci, Chang-Shou Lin, and Gabriella Tarantello.
\newblock Uniqueness and symmetry results for solutions of a mean field
  equation on {$\Bbb S^2$} via a new bubbling phenomenon.
\newblock {\em Comm. Pure Appl. Math.}, 64(12):1677--1730, 2011.

\bibitem{BartolucciTarantello-MR1917679}
Daniele Bartolucci and Gabriella Tarantello.
\newblock Liouville type equations with singular data and their applications to
  periodic multivortices for the electroweak theory.
\newblock {\em Comm. Math. Phys.}, 229:3--47, 2002.

\bibitem{Beckner-MR1230930}
William Beckner.
\newblock Sharp {S}obolev inequalities on the sphere and the
  {M}oser-{T}rudinger inequality.
\newblock {\em Ann. of Math. (2)}, 138(1):213--242, 1993.

\bibitem{BerlyandFuhrmannRybalko-MR3838409}
Leonid Berlyand, Jan Fuhrmann, and Volodymyr Rybalko.
\newblock Bifurcation of traveling waves in a {K}eller--{S}egel type free
  boundary model of cell motility.
\newblock {\em Commun. Math. Sci.}, 16(3):735--762, 2018.

\bibitem{CabreLuciaSanchon-MR2153523}
Xavier Cabre, Monica Lucia, and Maria Sanchon.
\newblock A mean field equation on a torus: one-dimensional symmetry of
  solutions.
\newblock {\em Comm. Partial Differential Equations}, 30(7--9):1315--1330,
  2005.

\bibitem{CaffarelliYang-MR1328259}
Luis~A. Caffarelli and Yisong~S. Yang.
\newblock Vortex condensation in the {C}hern--{S}imons {H}iggs model: an
  existence theorem.
\newblock {\em Comm. Math. Phys.}, 168(2):321--336, 1995.

\bibitem{CLMP1-MR1145596}
E.~Caglioti, P.-L. Lions, C.~Marchioro, and M.~Pulvirenti.
\newblock A special class of stationary flows for two-dimensional {E}uler
  equations: a statistical mechanics description.
\newblock {\em Comm. Math. Phys.}, 143(3):501--525, 1992.

\bibitem{CLMP2-MR1362165}
E.~Caglioti, P.-L. Lions, C.~Marchioro, and M.~Pulvirenti.
\newblock A special class of stationary flows for two-dimensional {E}uler
  equations: a statistical mechanics description. {II}.
\newblock {\em Comm. Math. Phys.}, 174(2):229--260, 1995.

\bibitem{Carlen-2025-StabilityLogHLS}
Eric~A. Carlen.
\newblock Stability for the logarithmic hardy–littlewood–sobolev
  inequality.
\newblock {\em Journal of Functional Analysis}, 295(2):Paper No. 110366, 2025.

\bibitem{ChanFuLin-MR1946446}
Hsin Chan, Ching-Chang Fu, and Chang-Shou Lin.
\newblock Non-topological multi-vortex solutions to the selfdual
  {C}hern--{S}imons--{H}iggs equation.
\newblock {\em Comm. Math. Phys.}, 231:189--221, 2002.

\bibitem{CLiu-MR1209959}
Kung~Ching Chang and Jia~Quan Liu.
\newblock On {N}irenberg's problem.
\newblock {\em Internat. J. Math.}, 4(1):35--58, 1993.

\bibitem{CCL-MR2055839}
Sun-Yung~A. Chang, Chiun-Chuan Chen, and Chang-Shou Lin.
\newblock Extremal functions for a mean field equation in two dimension.
\newblock In {\em Lectures on partial differential equations}, volume~2 of {\em
  New Stud. Adv. Math.}, pages 61--93. Int. Press, Somerville, MA, 2003.

\bibitem{CY2-MR925123}
Sun-Yung~A. Chang and Paul~C. Yang.
\newblock Conformal deformation of metrics on {$S^2$}.
\newblock {\em J. Differential Geom.}, 27(2):259--296, 1988.

\bibitem{ChangGui-CPA2023}
Sun Yung~Alice Chang and Changfeng Gui.
\newblock A sharp inequality on the exponentiation of functions on the sphere.
\newblock {\em Communications on Pure and Applied Mathematics},
  76(6):1303--1326, 2023.

\bibitem{CY1-MR908146}
Sun-Yung~Alice Chang and Paul~C. Yang.
\newblock Prescribing {G}aussian curvature on {$S^2$}.
\newblock {\em Acta Math.}, 159(3-4):215--259, 1987.

\bibitem{CK-MR1262195}
Sagun Chanillo and Michael Kiessling.
\newblock Rotational symmetry of solutions of some nonlinear problems in
  statistical mechanics and in geometry.
\newblock {\em Comm. Math. Phys.}, 160(2):217--238, 1994.

\bibitem{CGS-MR2944090}
Robin~Ming Chen, Yujin Guo, and Daniel Spirn.
\newblock Asymptotic behavior and symmetry of condensate solutions in
  electroweak theory.
\newblock {\em J. Anal. Math.}, 117:47--85, 2012.

\bibitem{CL1-MR1121147}
Wen~Xiong Chen and Congming Li.
\newblock Classification of solutions of some nonlinear elliptic equations.
\newblock {\em Duke Math. J.}, 63(3):615--622, 1991.

\bibitem{CL2-MR1338474}
Wen~Xiong Chen and Congming Li.
\newblock A necessary and sufficient condition for the {N}irenberg problem.
\newblock {\em Comm. Pure Appl. Math.}, 48(6):657--667, 1995.

\bibitem{DingJostLiWang-MR1634402}
Weiyue Ding, J{\"u}rgen Jost, Jiayu Li, and Guofang Wang.
\newblock An analysis of the two vortex case in the {C}hern--{S}imons {H}iggs
  model.
\newblock {\em Calc. Var. Partial Differential Equations}, 7(1):87--97, 1998.

\bibitem{DEJ-MR3377875}
Jean Dolbeault, Maria~J. Esteban, and Gaspard Jankowiak.
\newblock The {M}oser-{T}rudinger-{O}nofri inequality.
\newblock {\em Chin. Ann. Math. Ser. B}, 36(5):777--802, 2015.

\bibitem{DET-MR2509375}
Jean Dolbeault, Maria~J. Esteban, and Gabriella Tarantello.
\newblock Multiplicity results for the assigned {G}auss curvature problem in
  {$\Bbb R^2$}.
\newblock {\em Nonlinear Anal.}, 70(8):2870--2881, 2009.

\bibitem{GhoshKarmakar-arXiv2508.19930}
Monideep Ghosh and Debabrata Karmakar.
\newblock Quantitative stability for the conformally invariant chang-gui
  inequality on the exponentiation of functions on the sphere.
\newblock {\em arXiv preprint}, 2025.
\newblock Submitted 27 August 2025, 21 pages.

\bibitem{GL-MR2670931}
Nassif Ghoussoub and Chang-Shou Lin.
\newblock On the best constant in the {M}oser-{O}nofri-{A}ubin inequality.
\newblock {\em Comm. Math. Phys.}, 298(3):869--878, 2010.

\bibitem{GuiJevnikarMoradifam-MR3769320}
Changfeng Gui, Alessio Jevnikar, and Amir Moradifam.
\newblock Symmetry and uniqueness of solutions to some {L}iouville-type
  equations and systems.
\newblock {\em Comm. Partial Differential Equations}, 43(3):428--447, 2018.

\bibitem{GM}
Changfeng Gui and Amir Moradifam.
\newblock The sphere covering inequality and its applications.
\newblock {\em Invent. Math.}, 214(3):1169--1204, 2018.

\bibitem{GuiMoradifam-MR3762223}
Changfeng Gui and Amir Moradifam.
\newblock The sphere covering inequality and its applications.
\newblock {\em Invent. Math.}, 214:1169--1204, 2018.

\bibitem{GuiMoradifam-MR3741415}
Changfeng Gui and Amir Moradifam.
\newblock Uniqueness of solutions of mean field equations in $\mathbb{R}^2$.
\newblock {\em Proc. Amer. Math. Soc.}, 146(3):1231--1242, 2018.

\bibitem{GuiMoradifam-MR3927157}
Changfeng Gui and Amir Moradifam.
\newblock Symmetry of solutions of a mean field equation on flat tori.
\newblock {\em Int. Math. Res. Not.}, 2019:799--809, 2019.

\bibitem{GW-MR1760786}
Changfeng Gui and Juncheng Wei.
\newblock On a sharp {M}oser-{A}ubin-{O}nofri inequality for functions on
  {$S^2$} with symmetry.
\newblock {\em Pacific J. Math.}, 194(2):349--358, 2000.

\bibitem{K-MR1193342}
Michael K.-H. Kiessling.
\newblock Statistical mechanics of classical particles with logarithmic
  interactions.
\newblock {\em Comm. Pure Appl. Math.}, 46(1):27--56, 1993.

\bibitem{LeeLinTarantelloYang-MR3699389}
Youngae Lee, Chang-Shou Lin, Gabriella Tarantello, and Wen Yang.
\newblock Sharp estimates for solutions of mean field equations with collapsing
  singularity.
\newblock {\em Comm. Partial Differential Equations}, 42(10):1549--1597, 2017.

\bibitem{Lin1-MR1770683}
Chang-Shou Lin.
\newblock Uniqueness of solutions to the mean field equations for the spherical
  {O}nsager vortex.
\newblock {\em Arch. Ration. Mech. Anal.}, 153(2):153--176, 2000.

\bibitem{LinLucia-MR2263535}
Chang-Shou Lin and Monica Lucia.
\newblock Uniqueness of solutions for a mean field equation on torus.
\newblock {\em J. Differential Equations}, 229(1):172--185, 2006.

\bibitem{LinWang-MR2680439}
Chang-Shou Lin and Chin-Lung Wang.
\newblock Elliptic functions, {G}reen functions and the mean field equations on
  tori.
\newblock {\em Ann. of Math. (2)}, 172(2):911--954, 2010.

\bibitem{LinWang-MR3665624}
Chang-Shou Lin and Chin-Lung Wang.
\newblock Mean field equations, hyperelliptic curves, and modular forms: {I}.
\newblock {\em J. \'Ec. polytech. Math.}, 4:557--593, 2017.

\bibitem{LinWeiYe-MR2969270}
Chang-Shou Lin, Juncheng Wei, and Doudou Ye.
\newblock Classification and nondegeneracy of {$SU(n+1)$} {T}oda system.
\newblock {\em Invent. Math.}, 190(1):169--207, 2012.

\bibitem{Moser-MR0301504}
J.~Moser.
\newblock A sharp form of an inequality by {N}. {T}rudinger.
\newblock {\em Indiana Univ. Math. J.}, 20:1077--1092, 1970/71.

\bibitem{O-MR677001}
E.~Onofri.
\newblock On the positivity of the effective action in a theory of random
  surfaces.
\newblock {\em Comm. Math. Phys.}, 86(3):321--326, 1982.

\bibitem{OPS-MR960228}
B.~Osgood, R.~Phillips, and P.~Sarnak.
\newblock Extremals of determinants of {L}aplacians.
\newblock {\em J. Funct. Anal.}, 80(1):148--211, 1988.

\bibitem{PT-MR2876669}
Arkady Poliakovsky and Gabriella Tarantello.
\newblock On a planar {L}iouville-type problem in the study of selfgravitating
  strings.
\newblock {\em J. Differential Equations}, 252(5):3668--3693, 2012.

\bibitem{ShiSunTianWei-MRXXXXXXX}
Yuguang Shi, Jiacheng Sun, Gang Tian, and Dongyi Wei.
\newblock {\em Calc. Var. Partial Differential Equations}.

\bibitem{T-MR0216286}
Neil~S. Trudinger.
\newblock On imbeddings into {O}rlicz spaces and some applications.
\newblock {\em J. Math. Mech.}, 17:473--483, 1967.

\bibitem{WangWangYang-arXiv1804.03319}
Jun Wang, Zhi-An Wang, and Wen Yang.
\newblock Uniqueness and convergence on equilibria of the keller--segel system
  with subcritical mass.
\newblock arXiv:1804.03319 [math.AP], 2018.

\bibitem{Yang-MR1277473}
Yi~Song Yang.
\newblock Self-duality of the gauge field equations and the cosmological
  constant.
\newblock {\em Comm. Math. Phys.}, 162(3):481--498, 1994.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx, color}
\usepackage{hyperref}
\usepackage{epsfig}
\usepackage{epstopdf} 

\definecolor{darkgreen}{rgb}{0,0.55,0}

\newtheorem{proposition}{Proposition}[section]
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{remark}[theorem]{Remark}

\DeclareSymbolFont{AMSb}{U}{msb}{m}{n}
\DeclareMathSymbol{\N}{\mathbin}{AMSb}{"4E}
\DeclareMathSymbol{\Z}{\mathbin}{AMSb}{"5A}
\DeclareMathSymbol{\R}{\mathbin}{AMSb}{"52}
\DeclareMathSymbol{\Q}{\mathbin}{AMSb}{"51}
\DeclareMathSymbol{\I}{\mathbin}{AMSb}{"49}

\numberwithin{equation}{section}

\begin{document}

\title{\bf A Generalization of the Sphere Covering Inequality}
\author{{Changfeng Gui\footnote{Department of Mathematics, University of Macau, Macau SAR, P. R. China. E-mail address: changfenggui@um.edu.mo}
\qquad Amir Moradifam\footnote{Department of Mathematics, University of California, Riverside, California, USA. E-mail: moradifam@math.ucr.edu. }}}
\date{\today}

\smallbreak \maketitle

\begin{abstract}
The Sphere Covering Inequality was introduced in \cite{GM} (\emph{Invent. Math.}, 2018) as a sharp geometric inequality that provides a lower bound for the total area of two distinct surfaces of Gaussian curvature 1. These surfaces are assumed to be conformal to the Euclidean unit disk and share the same conformal factor along the boundary. In this paper, we establish a quantitative generalization that relaxes the boundary matching condition by allowing the conformal factors to differ by a constant \( c \ge 0 \) on the boundary. This refinement reveals a new stability-type structure underlying the inequality. Our results show that the Sphere Covering Inequality is stable with respect to perturbations in the boundary data and provide a precise quantitative description of how the total-area bound varies under such perturbations. The generalized inequality provides new analytic and geometric tools for the study of elliptic equations with exponential nonlinearities, conformal geometry, and related problems in mathematical physics.

\end{abstract}

\section{Introduction}

Second-order nonlinear elliptic equations with exponential nonlinearities form one of the most fundamental classes of equations in two-dimensional analysis. Equations of the form  
\begin{equation}\label{firstPDE}
\Delta u + e^{u} = f(x), \quad x \in \Omega \subset \mathbb{R}^2,
\end{equation}
arise in a wide variety of problems in mathematics, mathematical physics, and biology.  Equations of this form naturally arise in problems where curvature, conformal geometry, and analytic inequalities interact. In differential geometry, they describe surfaces with prescribed Gaussian curvature.  Analytically, they are closely related to the Moser–Trudinger type inequalities that capture the critical exponential growth characterizing the limiting case of Sobolev embeddings in dimension two. The interplay between geometric structures and sharp functional inequalities in this setting has led to deep insights into the nature of conformal invariance and the structure of extremal functions, see \cite{Aubin-MR0448404, Aubin2-MR534672, Beckner-MR1230930,  ChangGui-CPA2023, CL1-MR1121147, CL2-MR1338474, CLiu-MR1209959, CY1-MR908146, CY2-MR925123, DEJ-MR3377875, DET-MR2509375, GL-MR2670931, GW-MR1760786, Moser-MR0301504, O-MR677001, OPS-MR960228, T-MR0216286}.

In mathematical physics, such equations model a wide range of nonlinear phenomena. They govern the Chern–Simons self-dual vortex equations, Liouville and Toda systems, and mean-field equations arising in the statistical mechanics of two-dimensional turbulence and vortex dynamics. They also appear in the theory of self-gravitating cosmic strings in Einstein’s general relativity, where the exponential nonlinearity encodes the coupling between geometry and matter fields. Beyond these physical contexts, similar structures emerge in the theory of elliptic and hyperelliptic functions and in free-boundary problems such as models of cell motility and biological pattern formation; see \cite{BartolucciChenLinTarantello-MR2071339, BartolucciLin-MR2542689, BartolucciTarantello-MR1917679, BD-MR3355004, BerlyandFuhrmannRybalko-MR3838409, BLT-MR2838340, CabreLuciaSanchon-MR2153523, CaffarelliYang-MR1328259, CCL-MR2055839, CGS-MR2944090, ChanFuLin-MR1946446, CK-MR1262195, CLMP1-MR1145596, CLMP2-MR1362165, DingJostLiWang-MR1634402, K-MR1193342, Lin1-MR1770683, LinLucia-MR2263535, LinWang-MR2680439, LinWang-MR3665624, LinWeiYe-MR2969270, PT-MR2876669, Yang-MR1277473}.

Because of their critical nonlinearity and deep geometric origins, these equations provide a unifying framework that connects diverse areas of analysis, geometry, and mathematical physics. Their study has led to the discovery of sharp functional inequalities, refined classification results, and symmetry and uniqueness theorems for Liouville-type and mean-field equations. The interplay between the analytic structure of \eqref{firstPDE} and its geometric and physical interpretations continues to inspire new methods and results in modern nonlinear analysis.

A fundamental breakthrough in this area was the introduction of the \emph{Sphere Covering Inequality (SCI)} in~\cite{GM}, which asserts that the combined area of two distinct surfaces with Gaussian curvature equal to one, conformal to the Euclidean unit disk, and sharing the same conformal factor on the boundary, must be at least~$4\pi$. Geometrically, this means that the two surfaces together cover the entire unit sphere, up to a conformal rearrangement. Indeed, the following theorem was proved in \cite{GM}. \\

\begin{theorem}[The Sphere Covering Inequality \cite{GM}]\label{MainTheorem0}
Let $\Omega $ be a simply-connected  subset of $\R^2$ and assume $v_i \in C^2(\overline{\Omega})$, $i=1,2$ satisfy
\begin{equation}
\Delta v_i +e^{2v_i}=f_{i}(y),
\end{equation}
where $f_2 \geq f_1 \ge 0$  in $\Omega$.   If $v_2 \ge v_1,  v_2 \not \equiv v_1$ in $\omega$ and $v_2=v_1$ on $\partial \omega$  for some  piecewise
Liptschitz subdomain $\omega \subset \Omega$,  then
\begin{equation}
\int_{\omega} (e^{2v_1}+e^{2v_2})dy \geq 4\pi. 
\end{equation}
Moreover,  the equality only holds when $f_2 \equiv f_1 \equiv 0$  in $\omega$, and $(\omega, e^{2v_i}dy ) $, $i=1,2$ are isometric to  two complementary spherical caps on the standard unit sphere.  
\end{theorem}

The Sphere Covering Inequality was recently introduced in~\cite{GM} and has been applied to address several problems concerning symmetry and uniqueness of solutions to elliptic equations with exponential nonlinearities in dimension \( n = 2 \). In particular, it was employed to resolve a long-standing conjecture of Chang and Yang~\cite{CY2-MR925123} regarding the best constant in Moser--Trudinger type inequalities~\cite{GM}, as well as a conjecture of Bartolucci, Lin, and Tarantello~\cite{BLT-MR2838340} concerning mean field equations with singularities on~\(\mathbb{S}^2\) arising from self-dual gauge field theory. Moreover, it has led to a variety of further symmetry and uniqueness results for mean field equations, Onsager vortex models, the Sinh--Gordon and cosmic string equations, Toda systems, and rigidity phenomena related to the Hawking mass in general relativity (see \cite{BartolucciGuiJevnikarMoradifam-MR3961327, GuiJevnikarMoradifam-MR3769320, GuiMoradifam-MR3762223, GuiMoradifam-MR3927157, GuiMoradifam-MR3741415, LeeLinTarantelloYang-MR3699389, ShiSunTianWei-MRXXXXXXX, WangWangYang-arXiv1804.03319}). Beyond these specific applications, the Sphere Covering Inequality represents a central analytic principle in the treatment of nonlinear PDEs with critical exponential growth in two dimensions.

In this paper, we establish a generalization of the Sphere Covering Inequality that relaxes the boundary condition requiring two conformal metrics (or solutions) to coincide exactly on the boundary. Instead, we allow the solutions to differ by a constant \(c \ge 0\) on \(\partial \Omega\). This generalization broadens the applicability of the SCI to problems with nonidentical boundary data while preserving its geometric sharpness. Such a refinement provides new analytic tools for the study of elliptic PDEs with relaxed boundary conditions, symmetry-breaking phenomena, and geometric stability problems arising in conformal geometry and related fields.

Specifically, instead of assuming \(v_1 = v_2\) on \(\partial \Omega\), we permit a constant boundary gap:
\[
v_2 - v_1 = c \ge 0 \quad \text{on } \partial \Omega.
\]
Our main result shows that under these weaker boundary conditions one still obtains a quantitative lower bound, but with the constant \(4\pi\) replaced by a parameter depending on the boundary gap \(c\) relative to the maximal difference of the solutions inside \(\Omega\). Precisely, if \(\omega \subset \Omega\) is a Lipschitz subdomain on which \(v_2 \ge v_1\) and \(v_2 \not\equiv v_1\), then
\[
\int_{\omega}\big(e^{2v_1}+e^{2v_2}\big)\,dy \;\ge\; 4(1 - \frac{c}{M})\pi,
\]
where
\[
 M = \max_{\omega}(v_2 - v_1).
\]
When \(c = 0\), the result reduces to the classical Sphere Covering Inequality, and if $c=\frac{1}{2}\max\limits_{\omega} (v_2 - v_1)>0,$ then 
\[\int_{\omega} (e^{2v_1}+e^{2v_2})dy \geq 2 \pi.\]
Indeed we prove the following theorem which is the main result of this paper.  \\
\begin{theorem}[Generalized Sphere Covering Inequality]\label{GeneralSphereCoveringInequality}
Let \(\Omega \subset \mathbb{R}^2\) be simply connected, and assume \(v_i \in C^2(\overline{\Omega})\), \(i = 1, 2\), satisfy
\begin{equation}\label{mainpde}
\Delta v_i + e^{2v_i} = f_i(y),
\end{equation}
where \(f_2 \ge f_1 \ge 0\) in \(\Omega\). Suppose that \(v_2 \ge v_1+c\), \(v_2 \not\equiv v_1\) in a piecewise Lipschitz subdomain \(\omega \subset \Omega\), and that \(v_2 - v_1 = c \geq 0\) on \(\partial \omega\). Then
\begin{equation}\label{GeneralSCI}
\int_{\omega} (e^{2v_1} + e^{2v_2})\,dy \;\ge\; 4(1 - \frac{c}{M})\pi,
\end{equation}
where
\[
 M = \max_{\omega}(v_2 - v_1).
\]
Moreover, equality holds if and only if \(c = 0\), \(f_2 \equiv f_1 \equiv 0\) in \(\omega\), and the weighted surfaces \((\omega, e^{2v_i} dy)\), \(i = 1, 2\), are isometric to two complementary spherical caps on the standard unit sphere.
\end{theorem}

Conceptually, this result may be viewed as a stability-type extension of the Sphere Covering Inequality, quantifying how much geometric information is lost when the boundary agreement condition is relaxed. Quantitative stability results for sharp conformally invariant inequalities have recently been obtained in different settings: Carlen~\cite{Carlen-2025-StabilityLogHLS} established stability for the logarithmic Hardy--Littlewood--Sobolev inequality as well as for the Onofri inequality, and Ghosh and Karmakar~\cite{GhoshKarmakar-arXiv2508.19930} demonstrated quantitative stability for the Chang--Gui inequality on the sphere.

\begin{remark}
The assumption \( v_2 \ge v_1 + c \) in Theorem \ref{GeneralSphereCoveringInequality} can be replaced by the weaker condition 
\[
M=\max_{\omega}(v_2 - v_1) > c.
\]
Indeed, under this weaker hypothesis one may apply Theorem~\ref{GeneralSphereCoveringInequality} to  the set 
\[
\omega' :=\{ x \in \omega : v_2(x) - v_1(x) > c \},
\]
and conclude that 
\[
\int_{\omega} (e^{2v_1} + e^{2v_2})\,dy 
\;\ge\; 
\int_{\omega'} (e^{2v_1} + e^{2v_2})\,dy 
\;\ge\; 
4\!\left(1 - \frac{c}{M}\right)\!\pi.
\]
\end{remark}

\section{Auxiliary Results and Preliminaries}
In this section, we establish several preliminary results that will be used in the proof of the main theorem. These lemmas, though elementary in nature, play an essential role in clarifying the analytic framework and preparing the ground for the more delicate geometric arguments that follow.

\begin{lemma}\label{CalcLemma0}
For \( k \in [0,1] \), define
\[
f(x) = \frac{x^{2-k} - x^{k}}{x^{2} - 1} + k - 1.
\]
Then \( f(x) \le 0 \) for all \( x \in (1,\infty) \).  
In particular, \( f(x) < 0 \) for all \( x \in (1,\infty) \) if \( k \in (0,1) \).
\end{lemma}

{\bf Proof.}
First observe that \( f \equiv 0 \) for \( k \in \{0,1\} \).  
Assume \( k \in (0,1) \). Since
\[
\lim_{x \to 1} f(x) = 0,
\]
it suffices to show that \( f \) is decreasing on \( (1,\infty) \).  
A direct computation gives
\[
f'(x) = \frac{x\big( -k x^{2-k} + (2-k)x^{k} - (2-k)x^{-k} + kx^{k-2} \big)}{(x^{2}-1)^{2}},
\]
so \( f \) is decreasing if
\[
f_1(x) := -k x^{2-k} + (2-k)x^{k} - (2-k)x^{-k} + kx^{k-2} < 0
\]
for all \( x > 1 \).  
Since \( f_1(1) = 0 \), it is enough to verify that \( f_1'(x) < 0 \) on \( (1,\infty) \), where
\[
f_1'(x) = k(2-k)\big(-x^{1-k} + x^{k-1} + x^{-k-1} - x^{k-3}\big).
\]
Note that \( f_1'(x) < 0 \) if
\[
f_2(x) := -x^{1-k} + x^{k-1} + x^{-k-1} - x^{k-3}
\]
is decreasing on \( (1,\infty) \).  
Indeed,
\begin{align*}
f_2'(x)
&= -(1-k)x^{-k} - (1-k)x^{k-2} - (k+1)x^{-k-2} + (3-k)x^{k-4} \\
&< \big( -(1-k) - (1-k) - (k+1) + (3-k) \big)x^{k-4} = 0,
\end{align*}
and hence \( f_2 \) is strictly decreasing on \( (1,\infty) \).  
The proof is complete.\hfill \( \Box \) \\ \\

We shall also need the following elementary result.

\begin{lemma}\label{CalcLemma}
Let \( a, b, k \) be positive constants with \( b > a > 0 \) and \( k \in [0,1] \). 
Suppose that
\begin{equation}\label{CalcLemmaAssumption}
\frac{8 + b^2x^2}{8 + a^2x^2} = \left( \frac{b}{a} \right)^k
\end{equation}
holds. Then the inequality
\begin{equation}\label{CalclemmaInequality}
\frac{a^2x^2}{8 + a^2x^2} + \frac{b^2x^2}{8 + b^2x^2} \ge k
\end{equation}
is satisfied. Moreover, if \( k \in (0,1) \), then the inequality is strict.
\end{lemma}

{\bf Proof.}
From \eqref{CalcLemmaAssumption}, we obtain
\begin{equation}
x^2 = \frac{8\left( \left( \frac{b}{a} \right)^k - 1 \right)}{b^2 - \left( \frac{b}{a} \right)^k a^2},
\end{equation}
and hence
\begin{equation}
8 + a^2x^2 = \frac{8 \left( \left( \frac{b}{a} \right)^2 - 1 \right)}{\left( \frac{b}{a} \right)^2 - \left( \frac{b}{a} \right)^k}.
\end{equation}
Consequently,
\begin{equation}
\frac{8}{8 + a^2x^2} = \frac{\left( \frac{b}{a} \right)^2 - \left( \frac{b}{a} \right)^k}{\left( \frac{b}{a} \right)^2 - 1}.
\end{equation}
Therefore,
\begin{align*}
\frac{a^2x^2}{8 + a^2x^2} + \frac{b^2x^2}{8 + b^2x^2}
&= 2 - 8 \left( \frac{1}{8 + a^2x^2} + \frac{1}{8 + b^2x^2} \right) \\
&= 2 - \frac{8}{8 + a^2x^2} \left( 1 + \frac{1}{\frac{8 + b^2x^2}{8 + a^2x^2}} \right) \\
&= 2 - \frac{\left( \frac{b}{a} \right)^2 - \left( \frac{b}{a} \right)^k}{\left( \frac{b}{a} \right)^2 - 1} 
      \left( 1 + \frac{1}{\left( \frac{b}{a} \right)^k} \right) \\
&= 2 - \frac{\left( \frac{b}{a} \right)^2 - \left( \frac{b}{a} \right)^k}{\left( \frac{b}{a} \right)^2 - 1} 
      \frac{1 + \left( \frac{b}{a} \right)^k}{\left( \frac{b}{a} \right)^k} \\
&= 2 - \frac{\left( \frac{b}{a} \right)^{2 - k} - 1}{\left( \frac{b}{a} \right)^2 - 1}
      \left( 1 + \left( \frac{b}{a} \right)^k \right) \\
&= 2 - \frac{\left( \frac{b}{a} \right)^{2 - k} - 1 + \left( \frac{b}{a} \right)^2 - \left( \frac{b}{a} \right)^k}
             {\left( \frac{b}{a} \right)^2 - 1} \\
&= 1 - \frac{\left( \frac{b}{a} \right)^{2 - k} - \left( \frac{b}{a} \right)^k}
             {\left( \frac{b}{a} \right)^2 - 1}.
\end{align*}
Hence, inequality \eqref{CalclemmaInequality} holds if and only if
\begin{equation}
\frac{\left( \frac{b}{a} \right)^{2 - k} - \left( \frac{b}{a} \right)^k}
     {\left( \frac{b}{a} \right)^2 - 1} \le 1 - k.
\end{equation}
By Lemma~\ref{CalcLemma0}, the above inequality holds for \( \frac{b}{a} = x > 1 \).\hfill \( \Box \) \\ \\

For \( \lambda>0 \), define \( U_\lambda \) by 
\begin{equation}\label{ULambda}
U_{\lambda}:=-2\ln(1+\frac{\lambda^2 |y|^2 }{8})+2\ln(\lambda),
\end{equation}
which satisfies the equation 
\[
\Delta U_\lambda +e^{U_\lambda}=0.\\
\]
\\
We are now ready to prove the main result of this section.  \\

\begin{theorem}\label{RadialTheorem}
Let \( b > a > 0 \), and let \( U_a \) and \( U_b \) be defined as in~\eqref{ULambda}. 
Let \( k \in (0,1] \), and suppose that 
\begin{equation}\label{UAssumption}
U_b - U_a = 2(1 - k)\ln\!\left( \frac{b}{a} \right)
\quad \text{on } \partial B_R,
\end{equation}
where \( B_R \) is the ball of radius \( R \) centered at the origin. Then 
\begin{equation}
\int_{B_R} \!\big(e^{U_a} + e^{U_b}\big)\, dy \ge 8k\pi,
\end{equation}
and the inequality is strict if \( k \in (0,1) \).
\end{theorem}

{\bf Proof.}
From \eqref{UAssumption}, we have 
\[
-2 \ln\!\left( \frac{8 + b^2R^2}{8 + a^2R^2} \right) 
  + 2\ln\!\left( \frac{b}{a} \right) 
  = 2(1 - k)\ln\!\left( \frac{b}{a} \right),
\]
and hence 
\[
\frac{8 + b^2R^2}{8 + a^2R^2} = \left( \frac{b}{a} \right)^k.
\]
It then follows from Lemma~\ref{CalcLemma} that 
\begin{align*}
\int_{B_R} \!\big(e^{U_a} + e^{U_b}\big)\, dy
&= 8\pi \left( 
      \frac{a^2R^2}{8 + a^2R^2} 
    + \frac{b^2R^2}{8 + b^2R^2} 
   \right)
   \ge 8k\pi,
\end{align*}
with strict inequality when \( k \in (0,1) \).\hfill \( \Box \) \\ \\ 

\begin{remark}
It is important to note that Lemmas~\ref{CalcLemma0}, ~\ref{CalcLemma}, and Theorem \ref{RadialTheorem} are not valid for \( k > 1 \). Consequently, Theorem~\ref{GeneralSphereCoveringInequality} does not hold in the case \( c < 0 \).\\
\end{remark}

\section{Generalized Sphere Covering Inequality}

In this section, we present the proof of our main result, Theorem~\ref{GeneralSphereCoveringInequality}.  As a first step, we derive a comparison principle for radial solutions  involved, which will play a crucial role in our proof.

\begin{proposition}\label{LastEstimate}
Assume that $\psi \in C^{0, 1} (\overline{B_R})$ is a strictly decreasing, radial, Lipschitz function,  and  satisfies 
\begin{equation}\label{superSol}
\int_{\partial B_r} |\nabla \psi|ds  \le \int_{B_r}e^{\psi}dy
\end{equation}
{\it a.e.}  $r\in (0,R)$ and $\psi(0)=U_{b}(0)$  for some $b>0$. Then 
\begin{equation}\label{lastEstimate}
  \int_{B_R} e^{\psi} dy\geq   \int_{B_R} e^{U_{b}} dy.
\end{equation}
Moreover if the inequality in (\ref{superSol}) is strict at some point in $(0,R)$, then the inequality in (\ref{lastEstimate}) is also strict.
\end{proposition}
{\bf Proof.} Let $0<b'<b$.  We claim that 
\begin{equation}\label{bPrimeIneq}
\int_{B_R}e^{\psi}dy> \int_{B_R}e^{U_{b'}}dy. 
\end{equation}

Since $\psi(0)=U_{b}(0)>U_{b'}(0)$ and $\psi$ is continuous, we must have $\psi>U_{b'}$ in a neighborhood of the origin.  If  $\psi \geq U_{b'}$ in $B_R$, then \eqref{lastEstimate} follows immediately.  Define 
\[r_1=\sup\{s\in [0,R): \psi (r) > U_{b'}(r) \ \ \text{for all} \ \ r\in [0,s) \}.\]
Then $0<r_1< R$, $\psi(r_1)=U_{b'}(r_1)$, and 
\begin{equation}\label{larger}
\int_{B_{r_1}}e^{\psi(y)}dy>\int_{B_{r_1}}e^{U_{b'}(y)}dy. 
\end{equation}
On the other hand, by Bol's inequality we have 
\begin{eqnarray*}
\frac{1}{2}\left( \int_{B_{r_1}} e^\psi\right) \left( 8 \pi-\int_{B_{r_1}}e^\psi \right)&\leq &\left( \int_{\partial B_{r_1}}e^{\frac{\psi}{2}}\right)^2= \left( \int_{\partial B_{r_1}}e^{\frac{U_{b'}}{2}}\right)^2\\
&=&\frac{1}{2}\left( \int_{B_{r_1}} e^{U_{b'}}\right) \left( 8 \pi-\int_{B_{r_1 }}e^{U_{b'}} \right),
\end{eqnarray*}
and therefore $\int_{B_{r_1}}e^{\psi}>4\pi$. This follows from\eqref{larger} and the fact that the polynomial $\frac{1}{2}x(8\pi-x)$ is strictly  increasing on $(0, 4\pi)$.  Now define 
\[r_2=\sup \{s\in [0,R]: \int_{B_{s}}e^{\psi(y)}dy\geq\int_{B_{s}}e^{U_{b'}(y)}dy\},\]
and note that $r_2 \geq r_1.$ We claim that $r_2=R$. Suppose this is not the case.  Then there exists $r_3$ with $R \geq r_3>r_2$ such that 
\[\psi(r_3) < U_{b'}(r_3) \ \ \text{and}\ \  \int_{B_{r_3}}e^{\psi(y)}dy< \int_{B_{r_3}}e^{U_{b'}(y)}dy \]
By Bol's inequality we have 
\begin{eqnarray*}
\frac{1}{2}\left( \int_{B_{r_3}} e^\psi\right) \left( 8 \pi-\int_{B_{r_3}}e^\psi \right)&\leq &\left( \int_{\partial B_{r_3}}e^{\frac{\psi}{2}}\right)^2< \left( \int_{\partial B_{r_3}}e^{\frac{U_{b'}}{2}}\right)^2\\
&=&\frac{1}{2}\left( \int_{B_{r_3}} e^{U_{b'}}\right) \left( 8 \pi-\int_{B_{r_3 }}e^{U_{b'}} \right),
\end{eqnarray*}
this is a contradiction because $\frac{1}{2}x(8\pi-x)$ is increasing only for $x\leq 4\pi $ and $\int_{B_{r_3}}e^{\psi}dy>4\pi$.  Hence the inequality  \eqref{bPrimeIneq} holds for all $0<b'<b$.  The inequality \eqref{lastEstimate} follows from \eqref{bPrimeIneq} by letting $b' \rightarrow b$. The proof is now complete. \hfill $\Box$ \\ \\

Having established the necessary analytic framework and comparison principles, we are now ready to prove the central theorem of this work, the generalized Sphere Covering Inequality.  
It is worth noting that Theorem~\ref{GeneralSphereCoveringInequality} follows directly from the next result after a simple change of variable \( w = 2v \). \\

\begin{theorem}
Let \(\Omega \subset \mathbb{R}^2\) be simply connected, and assume \(v_i \in C^2(\overline{\Omega})\), \(i = 1, 2\), satisfy
\begin{equation}\label{mainpde}
\Delta v_i + e^{v_i} = f_i(y),
\end{equation}
where \(f_2 \ge f_1 \ge 0\) in \(\Omega\). Suppose that \(v_2 \ge v_1+c\), \(v_2 \not\equiv v_1\) in a piecewise Lipschitz subdomain \(\omega \subset \Omega\), and that \(v_2 - v_1 = c \geq 0\) on \(\partial \omega\). Then
\begin{equation}\label{GeneralSCI}
\int_{\omega} (e^{v_1} + e^{v_2})\,dy \;\ge\; 8(1 - \frac{c}{M})\pi,
\end{equation}
where
\[
 M = \max_{\omega}(v_2 - v_1).
\]
Moreover, equality holds if and only if \(c = 0\), \(f_2 \equiv f_1 \equiv 0\) in \(\omega\), and the weighted surfaces \((\omega, e^{2v_i} dy)\), \(i = 1, 2\), are isometric to two complementary spherical caps on the standard unit sphere.
\end{theorem}

{\bf Proof.} Suppose $v_1$ and $v_2$ satisfy the assumptions of the theorem. Then 
\[\Delta (v_2-v_1)+e^{v_2}-e^{v_1}=f_2-f_1\geq 0.\]
We may assume $\int_{\omega}e^{v_1}dy<4\pi$ and 
\begin{equation}\label{a-choice}
    \int_{\omega} e^{v_1}+e^{v_2}dy<8\pi-\epsilon
\end{equation}
for some $\epsilon >0$.  For any $a>0$,  there exists $R_a>0$ such that 
\begin{equation}
\int_{\omega}e^{v_1}dy=\int_{B_{R_a}}e^{U_{a}}dy= \frac{8\pi a^2R_{a}^2}{8+a^2R_{a}^2},
\end{equation}
where $U_{a}$ be define by \eqref{ULambda}. Observe that $a^2R_a^2=K=\text{constant}$, for any choice of $a>0$. Now let $b=a+M$ where $M:=\max\limits_{\omega} (v_2 - v_1)$. Hence, we can choose $a$ large enough such that 
\begin{eqnarray*}
    0<\int_{B_{R_a}}e^{U_b}dy-\int_{B_{R_a}}e^{U_a}dy&=& 8\pi \left[ \frac{(a+M)^2R_{a}^2}{8+(a+M)^2R_{a}^2}- \frac{a^2R_{a}^2}{8+a^2R_{a}^2}\right]\\
    &=& 8\pi \left[ \frac{(a+M)^2\frac{K}{a^2}}{8+(a+M)^2\frac{K}{a^2}}- \frac{K}{8+K}\right]<\epsilon.
\end{eqnarray*}

Fix $a$ large enough such that 
\begin{equation}\label{aa-choice}
    \int_{B_{R_a}}e^{U_b}dy-\int_{B_{R_a}}e^{U_a}dy <\epsilon,
\end{equation}
and for $t>c$ let $ \omega_t =  \{y \in \omega:    v_2(y) -v_1 (y)> t\}$. Let $\phi$ be the symmetrization of $v_2-v_1$ with respect to the measures $e^{v_1}dy$ and $e^{U_{a}}dy$.  More precisely,  define
\[\omega_t:=\{x\in \omega: v_2(x)-v_1(x)>t\},\]
and define $\omega^*_t$ to be the ball centered at the  origin in $\R^2$ such that 
\[\int_{\omega^*_t}e^{U_{a}}dy=\int_{\omega_t}e^{v_1}dy:=\beta(t).\]

Then $\beta(t) $ is  a right-continuous function,  and $\phi: \omega \rightarrow \R$ defined by $\phi(y):=\sup \{t\in \R: y\in \omega^*_t\}$ provides an equimeasurable rearrangement of $v_2-v_1$ with respect to the measure $e^{v_1}dy$ and $e^{U_{a}}dy$, i.e. 
\begin{equation}\label{rearrang}
\int_{\{\phi>t\}}e^{U_{a}}dy=\int_{\{v_2-v_1>t\}}e^{v_1}dy, \ \ \forall  t>\min_{y\in \overline \omega}\phi,
\end{equation}
where $\omega^*$  a ball of radius $R_a<\infty$ and centered at the origin with $\int_{\omega^*}e^{U_{a}}dy=\int_{\omega}e^{v_1}dy$.

By Propositions 2.2 and 2.3 in \cite{GM},   Green's formula,   and Cavalieri's principle we have  
\begin{eqnarray} \label{firstIneq}
\int_{\{\phi=t\}}|\nabla \phi|ds &\leq &  \int_{\{v_2-v_1=t\}\cap \omega}|\nabla (v_2-v_1)|  ds    \\
&\leq&\int_{\omega_t} \bigl(e^{v_2}-e^{v_1} \bigr) dy   \quad \hbox{ (by Green's formula and equation \eqref{mainpde} ) }  \nonumber \\
&=& \int_{\omega_t}e^{v_2-v_1} e^{v_1}dy -\int_{\omega_t}e^{v_1}  dy \nonumber \\
&=& \int_{\{\phi>t\}}e^{\phi}e^{U_{a }}dy- \int_{\{\phi>t\}}e^{U_{a}} dy \nonumber \\
&&  \quad \hbox{ (by rearrangement and Cavalieri's principle) }  \nonumber \\
&=& \int_{\{\phi>t\}}e^{U_{a}+\phi}dy-\int_{\{\phi=t\}} |\nabla U_{a}|ds,  \quad  \hbox{for  {\it a.e.} }  t>c.\nonumber
\end{eqnarray} 
  Hence
\begin{equation}\label{Sub}
\int_{\{\phi=t\}} |\nabla (U_{a}+\phi)| ds\le \int_{\{\phi>t\} } e^{(U_{a}+\phi)}dy,  \quad  \hbox{for  {\it a.e.} }  t>c.
\end{equation}
  Since  $ \phi \geq c$ is decreasing in $r$,     $\psi:= U_{a}+\phi$ is a strictly decreasing radial function, and 
\begin{equation}\label{supersolution}
\int_{\partial B_r} |\nabla \psi|ds \le  \int_{B_r} e^{\psi} dy, \quad {\it a.e.}  \quad r \in (0, R_a), 
\end{equation}
by Proposition 2.3 in \cite{GM} and the above inequality we know that  $\psi$ belongs to  $C^{0, 1} (B_{R_a})$.  It follows from Proposition \ref{LastEstimate} that 
\[\int_{B_{R_a}}e^{\psi}dy =\int_{B_{R_a}}e^{U_{a}+\phi}dy \geq 
\int_{B_{R_a}}e^{U_{b}}dy.\]

Next we claim that $\psi(R_a) \geq U_{b} (R_a)$. Suppose that this is not the case. Since, by Proposition \ref{LastEstimate},
\[\int_{B_r}e^{\psi}dy \geq \int_{B_r}e^{U_b}dy, \ \ \forall r\in (0,R_a),\]
there must exist $R_0 \in (0,R_a)$ such that $\psi (R_0)=U_b(R_0)$, and  $\int_{B_{R_0}}e^{\psi}dy > \int_{B_{R_0}}e^{U_b}dy$. Thus it follows from the Sphere Covering Inequality that  
\[\int_{B_{R_0}}e^{\psi}+e^{U_b}dx \geq 8\pi.\]
Thus it follows from \eqref{aa-choice} that
\begin{eqnarray*}
    \int_{\omega}e^{v_1}+e^{v_2}dy &\geq& \int_{B_{R_0}}e^{U_a}+e^{\psi}dx\\
    &\geq & \int_{B_{R_0}}e^{U_b}+e^{\psi}dx-\epsilon \\
    &\geq & 8\pi -\epsilon. 
\end{eqnarray*}
This is a contradiction in view of \eqref{a-choice}, and hence $$\psi(R_a) \geq U_{b} (R_a)$$.  

Now we have
\begin{eqnarray*}
\int_{\Omega}(e^{v_1}+e^{v_2})dy&=& \int_{B_{R_a}}(e^{U_{a}}+e^{\psi})dy \\ &\geq& \int_{B_{R_a}}(e^{U_{a}}+e^{U_{b}})dy\geq 8\pi \left[  1-\frac{U_b(R_a)-U_a({R_a})}{M}\right] \\ 
&\geq& 8\pi \left[  1-\frac{\psi(R_a)-U_a({R_a})}{M}\right]\\
&= &8\pi \left[  1-\frac{(v_2-v_1)|_{\partial \omega}}{M}\right] = 8(1 - \frac{c}{M})\pi.
\end{eqnarray*}
Moreover,  if  the equality holds,  then  $k=1$, and the inequality \eqref{GeneralSCI} simplifies to the Sphere Covering Inequality \cite{GM} for which the equality case is established. \hfill $\Box$ \\

$\mathbf{Acknowledgement}.$ The first author is  partially supported by University of Macau research grants (No. CPG2024-00016-FST, CPG2025-00032-FST, SRG2023-00011-FST and MYRGGRG2023-00139-FST-UMDF), UMDF Professorial Fellowship of Mathematics, Macao SAR FDCT 0003/2023/RIA1 and Macao SAR FDCT 0024/2023/RIB1. The second author is supported in part by NSF grant DMS-1715850.

\bibliographystyle{plain}
\bibliography{GeneralizedSphereCoveringInequality.bib}

\end{document}